OpenStudy (anonymous):

What value of y makes the equation true? 3.4 + y = –6.85 A. –10.25 <----- B. –3.25 C. 3.25 D. 10.25

OpenStudy (phi):

to "solve for y" add -3.4 to both sides, and simplify 3.4 + y = –6.85 3.4 - 3.4 + y = -6.85 -3.4 0+y= -10.25 y= -10.25